A man steps off an overhanging cliff and falls 9m into a lake. How long does it take him to hit the water? How fast is he falling when he reaches the lake’s surface?
1
Expert's answer
2019-10-31T11:20:47-0400

The equations of motion of a man are "y(t) = h - \frac{g t^2}{2}" and "v(t) = gt", where "h" is the height of the cliff, "g = 9.81 \frac{m}{s^2}"- gravitational acceleration. Equating "y(t)" to zero, and solving for "t", obtain "0 = h - \frac{g t^2}{2} \Rightarrow t = \sqrt{\frac{2 h}{g}}" - this is the time it takes to fall down. Calculating speed at that moment, obtain "v = g \sqrt{\frac{2 h}{g}} = \sqrt{2 g h}".

Calculating for "h = 9 m", obtain "t \approx 1.35 s", "v \approx 13.29 \frac{m}{s}".